22west:

Melissa needs 3.87 feet of string to use as a border for her poster. Melissa finds a piece of string that is 4.25 feet long. How much string will Melissa have left over?

22west:

@Ultrilliam @Allison

22west:

@lexxii

Vocaloid:

subtract total string - string used for the border

Vocaloid:

4.25 - 3.87 = ?

22west:

0.38?

Vocaloid:

good, that's it (make sure to include units, so 0.38 feet)

22west:

ok

22west:

Marcus is making cookies. The recipe says 2 1/4 cups of flour are needed for each batch. Marcus wants to make 3 batches of cookies. How many cups of flour will Marcus use? 3/4 cups 1 1/3 cups 5 1/4 cups 6 3/4 cups / = fraction

22west:

@Vocaloid

Vocaloid:

multiply cups per batch * number of batches

Vocaloid:

(2 and 1/4)*3 = ?

22west:

idk

Vocaloid:

split 2 and 1/4 into (2+1/4) 2*3 + 2*(1/4) = ?

22west:

6.5?

Vocaloid:

whoops made a typo

Vocaloid:

2*3 + 3*(1/4) = ?

22west:

6.75?

Vocaloid:

good, now convert that to a fraction

22west:

6 3/4

Vocaloid:

good so that's your answer.

22west:

i have more

22west:

The Harper family is traveling to visit friends. The distance to the friend’s house is 276.45 miles. It takes the family 5.82 hours to complete the trip. Rate is found by dividing distance by time. What is the rate for the family’s trip in miles per hour?

Vocaloid:

"Rate is found by dividing distance by time." so divide your distance (276.45) by the time (5.82)

22west:

47.5

Vocaloid:

good, that's it

22west:

Ian wants to run 4 1/2 miles to get ready for his upcoming race. He has set up 3 markers the same distance apart to show how far he has run. How far apart are the markers?

Vocaloid:

so if you put 3 markers down, you actually divide the distance by 4

Vocaloid:

so divide 4 1/2 by 4

22west:

1.125

Vocaloid:

good now convert that back into a fraction

22west:

1 1 --- 8

22west:

1 1/8

Vocaloid:

good, that's it

22west:

the answers are 3/2, 15/2, and 27/2

Vocaloid:

oh, perhaps they wanted a division by 3 instead :S

22west:

ima guess

22west:

i said 3/2

Vocaloid:

good, 3/2 is it

22west:

The water level at Lake Penny changed −1 1/10 in. this year. Next year, it is predicted to change another −1 2/5 in. What is the total change that will likely occur in 2 years?

Vocaloid:

add the two numbers together

Vocaloid:

(-1 1/10) + (-1 2/5) = ?

22west:

-2.5

Vocaloid:

good now convert that back into a fraction

Vocaloid:

as a general rule if they give you fractions the answer should also be in fractions

22west:

\[-2\frac{ 1 }{ 2 }\]

Vocaloid:

good that's it.

22west:

100% BOI

22west:

brb